OpenStudy (anonymous):

Suppose you invest $7500 at an annual interest rate of 4.2% compounded continuously. How much will you have in the account after 2 years? Round the solution to the nearest dollar.

OpenStudy (anonymous):

Interest=principle x rate x time (in years) Interest=7500 (.042)(2) Interest=630 7500+630=8130 Interest=500 (.071)(4) Interest=142 500+142=642

OpenStudy (anonymous):

\[\large 7500\times e^{.042\times 2}\] and calculator